# LLM Inference Unveiled: Survey and Roofline Model Insights

### v5

@@ 4.2.1 Speculative Decoding
Knowledge Distillation and Self-Speculative Decoding Another way to improve the acceptance rate is to improve the small draft model’s ability to align with the LLM’s generation distribution, which can be done through finetuning the small models on corpus generated by the large models with knowledge distillation. Zhou et al., 2023b establishes a mathematical connection between the acceptance rate and natural divergence between the small model and the LLM: minimizing the divergence is maximizing the acceptance rate. The paper also studies a range of different knowledge distillation losses and shows that adding knowledge distillation brings consistent 10-45% improvement in latency speedup. However, the paper generally finds that the optimal knowledge distillation loss choices vary model by model and should be tuned as a hyperparameter. Liu et al., 2023a also shows that knowledge distillation boosts the small model training. Besides, the paper brings speculative decoding to the cloud online learning settings. LLM inference is memory-bottlenecked, which means that there is always a surplus in computation resources. The compute can be used to train a draft model continously on server, which brings two benefits: Continuously training with knowledge distillation boosts its acceptance rate and, thus, reduces the LLM inference latency; 2) serving input is constantly shifting in domains, and continuous training helps the draft models maintain the strong performance in different domains. Zhang et al., [2023] avoids storing a separate draft model by selectively sampling a smaller draft model from the large model itself. Before deployment, the paper utilizes a Bayesian optimization method to search for a draft model by skipping intermediate layers within the pretrained large model. Besides, it proposes an adaptive threshold selection technique tailored for the decoding of the draft model sampled from the large models.
@@ 4.2.2 Parallel Decoding
Non-Autoregressive Transformers For Machine Translation tasks that require autoregressive decoding of the sequence-to-sequence model, Non-Autoregressive Transformers (NAT) has been proposed to iteratively decode all of the output tokens together, as shown in Figure 13 (d). NAT has been relatively well-explored (Gu et al., [2017], Wang et al., [2019], Li et al., [2019], Sun et al., 2019b , Wei et al., [2019], Shao et al., [2020], Lee et al., [2018], Ghazvininejad et al., [2019], Guo et al., [2020], Gu and Kong, [2020], Savinov et al., [2021]), and we point the readers to the following survey paper that covers specifically NAT models Xiao et al., 2023b for an in-depth review and analysis on the subject. Coarsely, the speedup of text decoding comes from making a single forward pass of the decoder output more than one token. The input sequence is first fed into the encoder, which outputs the hidden states that extract the input semantics. The output hidden states of the encoder are then used as the condition for the decoder pass. To speed up the text generation, the decoder side relaxes the autoregressive constraints and takes a sequence full of dummy tokens [pad] as the input to start the iterative parallel decoding process. During each iteration, based on the condition set by the encoder output hidden states, some tokens can be confidently predicted, which are unmasked. The sequence is mixed with unmasked decoded tokens and the remaining masked tokens are fed to the decoder again until every token is decoded. The length of the sequence fed into the decoder, or fertility, is usually learned either inside the encoder as a special [CLS] token or by a specialized fertility predictor between the encoder and the decoder. More recently, Savinov et al., [2021] treats the decoder as a diffusion model and trains it to denoise the noisy initial sequence based on the conditions given. However, because of the requirement to use encoder hidden states as the condition for parallel decoding, NAT methods face natural difficulties in extending directly to decoder-only architectures.
@@ 6.2 Processing in Memory
  • Inadequate Inter-PIM Communication. In addition to the constraints imposed by computation power and capacity, another limitation of DRAM-PIMs is their suboptimal inter-PIM communication capability. Given that there are distributed computing units located near each DRAM bank, data aggregation and computation synchronization among these units are unavoidable. However, DRAM-PIMs lack robust interconnects Zhou et al., 2023c , Jonatan et al., [2024], and they typically depend on the host CPU/GPU for data exchange between PIM units. This reliance can lead to inefficiencies in the system. Therefore, to enhance LLM inference, future iterations of DRAM-PIMs should aim to improve their inter-PIM communication abilities.
@@ 6.3 New Data Format
On the other hand, researchers are delving into the refinement of low-precision quantization formats to augment the adaptability of data representation whilst preserving hardware efficiency. One line of works propose to explore new encoding schemes within single value representation. Contrary to INT and FP numbers, which utilize fixed-length sub-fields for encoding distinct pieces of information, such as exponents and mantissas, the newly-proposed rule-based quantization formats enable dynamic adjustment of sub-field bit-widths. ALPS [Langroudi et al., , 2021] proposes a generalized posit format along with a new adaptive quantization algorithm to optimally represent the dynamic range and distribution of DNN parameters. ANT [Guo et al., , 2022] proposes a new data format called flint with leading-1 encoding for exponent fields. Dybit [Zhou et al., 2023a, ] proposes to separate exponent and mantissa fields with the first encountered 0 as delimiter. The flexibility of these variable-length data formats presents the chance to compromise between range and precision more effectively and allows for customization to more closely align with the distribution of LLMs’ weights and activations.
@@ 7.1 Reliability
Hallucination primarily refers to cases where the outputs of LLMs do not align with real-world knowledge, often generating content that is either factually incorrect or nonsensical Huang et al., [2023]. Safety alignment focuses on the model’s inherent ability to autonomously recognize and refuse to respond to harmful inquiries, thereby safeguarding against generating inappropriate or dangerous content Ouyang et al., [2022]. Reliability pertains to the model’s stability when confronted with unconventional data in long-tail scenarios, such as interference from adversarial examples or with decision shortcuts Geirhos et al., [2020]. In the following parts, we will delve into how different compression methods impact these three crucial aspects of model performance.
@@ 7.4.1 Large Multimodal Models (LMMs)
Large Multimodal Models (LMMs), particularly Visual Language Models (VLMs), have emerged as a promising avenue for creating general-purpose assistants, showcasing significant enhancements in perception and reasoning capabilities. These models leverage LLMs as their cognitive core, enriching multimodal (MM) tasks with robust language generation, zero-shot transfer capabilities, and In-Context Learning. Foundation models across different modalities provide high-quality representations. A critical challenge for LMMs is integrating LLMs with models from other modalities to facilitate collaborative inference effectively. The primary focus has been on improving modality alignment and aligning with human intentions through a MM Pre-Training + MM Instruction-Tuning pipeline. Two surveys, [Yin et al., 2023b, , Zhang et al., , 2024], delve into LMMs in detail.

### v6

@@ 4.1.3 Mixture-of-Expert Models
4.1.4 Roofline Model Analysis for Dynamic Parameter Reducing
The Minimum Parameter Used Per Token Decoded methods simultaneously decrease computational and memory access overhead. From the viewpoint of roofline model, these methods result in small changes to the arithmetic intensity of each operator and the type of bound.
For Early Exiting or Layer Skipping methods, entire Transformer layers are skipped, leading to a proportional reduction in overall computation, memory access, and inference time. In other words, the inference time decreases proportionally to the number of layers skipped in the network. However, for methods like Contextual Sparsity and Mixture of Experts, the arithmetic intensity varies across different operations. Consequently, dynamically choose to activate these layers leads to varying reductions in computation and memory access, resulting in different impacts on the overall inference time.
@@ 4.2.1 Speculative Decoding
Knowledge Distillation and Self-Speculative Decoding Another way to improve the acceptance rate is to improve the small draft model’s ability to align with the LLM’s generation distribution, which can be done through finetuning the small models on corpus generated by the large models with knowledge distillation. Zhou et al., 2023c establishes a mathematical connection between the acceptance rate and natural divergence between the small model and the LLM: minimizing the divergence is maximizing the acceptance rate. The paper also studies a range of different knowledge distillation losses and shows that adding knowledge distillation brings consistent 10-45% improvement in latency speedup. However, the paper generally finds that the optimal knowledge distillation loss choices vary model by model and should be tuned as a hyperparameter. Liu et al., 2023a also shows that knowledge distillation boosts the small model training. Besides, the paper brings speculative decoding to the cloud online learning settings. LLM inference is memory-bottlenecked, which means that there is always a surplus in computation resources. The compute can be used to train a draft model continously on server, which brings two benefits: Continuously training with knowledge distillation boosts its acceptance rate and, thus, reduces the LLM inference latency; 2) serving input is constantly shifting in domains, and continuous training helps the draft models maintain the strong performance in different domains. Zhang et al., [2023] avoids storing a separate draft model by selectively sampling a smaller draft model from the large model itself. Before deployment, the paper utilizes a Bayesian optimization method to search for a draft model by skipping intermediate layers within the pretrained large model. Besides, it proposes an adaptive threshold selection technique tailored for the decoding of the draft model sampled from the large models.
@@ 4.2.2 Parallel Decoding
Non-Autoregressive Transformers For Machine Translation tasks that require autoregressive decoding of the sequence-to-sequence model, Non-Autoregressive Transformers (NAT) has been proposed to iteratively decode all of the output tokens together, as shown in Figure 13 (d). NAT has been relatively well-explored (Gu et al., [2017], Wang et al., [2019], Li et al., [2019], Sun et al., 2019b , Wei et al., [2019], Shao et al., [2020], Lee et al., [2018], Ghazvininejad et al., [2019], Guo et al., [2020], Gu and Kong, [2020], Savinov et al., [2021]), and we point the readers to the following survey paper that covers specifically NAT models Xiao et al., 2023c for an in-depth review and analysis on the subject. Coarsely, the speedup of text decoding comes from making a single forward pass of the decoder output more than one token. The input sequence is first fed into the encoder, which outputs the hidden states that extract the input semantics. The output hidden states of the encoder are then used as the condition for the decoder pass. To speed up the text generation, the decoder side relaxes the autoregressive constraints and takes a sequence full of dummy tokens [pad] as the input to start the iterative parallel decoding process. During each iteration, based on the condition set by the encoder output hidden states, some tokens can be confidently predicted, which are unmasked. The sequence is mixed with unmasked decoded tokens and the remaining masked tokens are fed to the decoder again until every token is decoded. The length of the sequence fed into the decoder, or fertility, is usually learned either inside the encoder as a special [CLS] token or by a specialized fertility predictor between the encoder and the decoder. More recently, Savinov et al., [2021] treats the decoder as a diffusion model and trains it to denoise the noisy initial sequence based on the conditions given. However, because of the requirement to use encoder hidden states as the condition for parallel decoding, NAT methods face natural difficulties in extending directly to decoder-only architectures.
@@ 6.2 Processing in Memory
  • Inadequate Inter-PIM Communication. In addition to the constraints imposed by computation power and capacity, another limitation of DRAM-PIMs is their suboptimal inter-PIM communication capability. Given that there are distributed computing units located near each DRAM bank, data aggregation and computation synchronization among these units are unavoidable. However, DRAM-PIMs lack robust interconnects Zhou et al., 2023d , Jonatan et al., [2024], and they typically depend on the host CPU/GPU for data exchange between PIM units. This reliance can lead to inefficiencies in the system. Therefore, to enhance LLM inference, future iterations of DRAM-PIMs should aim to improve their inter-PIM communication abilities.
@@ 6.3 New Data Format
On the other hand, researchers are delving into the refinement of low-precision quantization formats to augment the adaptability of data representation whilst preserving hardware efficiency. One line of works propose to explore new encoding schemes within single value representation. Contrary to INT and FP numbers, which utilize fixed-length sub-fields for encoding distinct pieces of information, such as exponents and mantissas, the newly-proposed rule-based quantization formats enable dynamic adjustment of sub-field bit-widths. ALPS [Langroudi et al., , 2021] proposes a generalized posit format along with a new adaptive quantization algorithm to optimally represent the dynamic range and distribution of DNN parameters. ANT [Guo et al., 2022a, ] proposes a new data format called flint with leading-1 encoding for exponent fields. Dybit [Zhou et al., 2023a, ] proposes to separate exponent and mantissa fields with the first encountered 0 as delimiter. The flexibility of these variable-length data formats presents the chance to compromise between range and precision more effectively and allows for customization to more closely align with the distribution of LLMs’ weights and activations.
@@ 7.1 Reliability
Hallucination primarily refers to cases where the outputs of LLMs do not align with real-world knowledge, often generating content that is either factually incorrect or nonsensical Huang et al., 2023a . Safety alignment focuses on the model’s inherent ability to autonomously recognize and refuse to respond to harmful inquiries, thereby safeguarding against generating inappropriate or dangerous content Ouyang et al., [2022]. Reliability pertains to the model’s stability when confronted with unconventional data in long-tail scenarios, such as interference from adversarial examples or with decision shortcuts Geirhos et al., [2020]. In the following parts, we will delve into how different compression methods impact these three crucial aspects of model performance.
@@ 7.4.1 Large Multimodal Models (LMMs)
Large Multimodal Models (LMMs), particularly Visual Language Models (VLMs), have emerged as a promising avenue for creating general-purpose assistants, showcasing significant enhancements in perception and reasoning capabilities. These models leverage LLMs as their cognitive core, enriching multimodal (MM) tasks with robust language generation, zero-shot transfer capabilities, and In-Context Learning. Foundation models across different modalities provide high-quality representations. A critical challenge for LMMs is integrating LLMs with models from other modalities to facilitate collaborative inference effectively. The primary focus has been on improving modality alignment and aligning with human intentions through a MM Pre-Training + MM Instruction-Tuning pipeline. Two surveys, [Yin et al., 2023b, , Zhang et al., 2024a, ], delve into LMMs in detail.
@@ 7.4.2 Efficient LMMs
7.5 Long Context Modeling
When used for tasks like Chatbot or document summarization tools, Large Language Models’ long context language modeling and reasoning capabilities are challenged. The model, however, are usually trained on the general pretraining corpus that usually comprises text snippets, not long enough to serve as high-quality training examples for LLM to learn. To alleviate the insufficient long context capabilities of the pre-trained models, abundant works have attempted to attach the problem from different angles. For this section of the discussion, we mainly focus on alternative attention mechanisms in 7.5.1, cache compression and context retrieval, and position encoding modifications. For more detailed studies and a more comprehensive review of the issue of LLM long-context modeling, we refer interested readers to the recent survey specifically on this topic Huang et al., 2023b .
7.5.1 Alternative Attention Design
Lying in the core of transformer architecture is the self-attention mechanism. For decoder-only model inference with KV cache, if the past context is long, attending to all past keys computing with past values incurs both computation and memory bottleneck. Prior works found that not all past tokens need to be attended to preserve the model inference performance. Landmark Attention (Mohtashami and Jaggi, [2023]) introduced the special landmark attention into the sequence to summarize the following block of the tokens’ information. The new query would attend to the landmark tokens first to determine whether the following tokens inside the block are needed for predicting the next word, thus reducing the attention computation while maintaining the random-access nature of attention. Earlier, Funnel-Transformer (Dai et al., [2020]) also approaches the same goal of conducting attention in the block level. Differently, they introduce a down-sampling during the encoder portion and an up-sampling method during the decoder portion. The reduced FLOPs allow them to build a deeper and wider model that outperforms the original model with the same compute budget. On the other hand, Longformer (Beltagy et al., [2020]) gives an early attempt to combine sliding window attention and global attention, in which they only allow a few pre-defined tokens to attend to all tokens in the sequence, while other tokens do sliding window attention and also attend to these selected global attending tokens. Concurrently, ETC (Ainslie et al., [2020]) introduces a 2-level hierarchy to the input so that the plain long input tokens can do sliding window attention, while a global input of auxiliary tokens extracted and downsampled from the original input can do normal attention. The plain inputs are allowed to attend to the global inputs, thus acquiring the global context information. Similarly, LongT5 (Guo et al., 2022b ) proposes an even simpler method to directly downsampled the long and prior context using mean pooling with chunksize of 16, the downsampled keys and values are directly appended in front of the rest of the input sequence. Then the rest of the model can do sliding window attention with the additional attention to the downsampled context summary tokens at the front. Under the LLM era when pretraining becomes prohibitively costly, StreamingLLM (Xiao et al., 2023b ) and LM Infinite (Han et al., [2023]) concurrently propose a sink plus sliding window attention pattern to LLM as a plug-in tool to boost the LLM’s long context ability. StreamingLLM in particular points out that due to the mechanism of softmax operation in the transformer, the beginning tokens of the input are crucial to maintaining the self-attention performance. The modified attention mask is shown to achieve strong results in long-context language modeling without additional model finetuning. Besides, $H_{2}O$ (Zhang et al., 2024b ) reduces the self-attention computational complexity by only attending to tokens of interest out from the prior prefilled input context. To do that, they build an empirical oracle for selecting tokens. The method is shown to benefit LLM in long-context regime.
7.5.2 Recurrence and Retrieval
Transformer-XL Dai et al., [2019] proposes to introduce the segment-level recurrence structure to the Language Model to boost the current language model in long-context capabilities. The method stores the last layer output of one previous segment to append to the current layer, thus drastically increasing the dependency distance of the model. Segatron (Bai et al., [2021]) and Compressive Transformer (Rae et al., [2019]) extend upon the previous idea. Segatron boosts the segment-level recurrence from the position embeddings on the token level, sentence level, and beyond through segment-aware mechanisms. Compressive Transformer proposes a second-level compressed memory FIFO queue so that the past segment context should not be discarded but compressed by their customized function and stored in the queue to elongate the length of context dependency. In the LLM era, Dynamic Memory Compression Nawrot et al., [2024] also follows the recurrence idea with compressed context to dynamically decide how to compress the previous context information, thus reducing the sequence length of the attention while preserving the distant information. Besides, other than segment-level, Fan et al., [2020] studies the retrospective recurrence. Memorizing Transformer (Wu et al., [2022]) and Memformer (Wu et al., [2020]) combines retrieval and local cache with a forgetting mechanism.
On the other hand, following on trend that not every past token is needed for the current token generation next, the past KV cache can be placed physically further in distance, i.e. secondary memory, so that when only needed, the specific key-value pair will be retrieved. Thus, another way to boost the LLM is through Retrieval-Augmented Generation (RAG). RAG is a heated topic in its own light and contains abundant techniques and past works. Due to the limited scope of our paper, we kindly refer to the following comprehensive surveys on RAG for interested readers: Gao et al., [2023] and Zhao et al., [2024]. To address the particular motivation underlined long-context capabilities brought up at the beginning of this section, LangChain Pandya and Holia, [2023] is the popular way to mitigate the Chatbot’s past long conversation through retrieval. LangChain is an open-source tool that specializes in using LLM to compute embedding for user-input long documents and files, so that later according to the user’s prompt, top-relevant contents will be retrieved through the cosine-similarity metrics. Besides, there are fast-growing other relevant works(Borgeaud et al., [2021], Bertsch et al., [2024], Zhong et al., [2023], Zhou et al., 2023b , Kynoch et al., [2023], Modarressi et al., [2023], Guu et al., [2020], Wang et al., 2024b ) under the retrieval for long-context settings.
7.5.3 Maneuvering Position Encodings
During pretraining, the position encoding of the transformer hasn’t seen an input sequence length longer than a fixed limit. Also, because the position encoding is usually based on a triangular function, the vanilla transformer is unable to extrapolate to unfamiliar longer sequence lengths. Earlier techniques add attention bias to the attention map before the softmax operation. ALiBi (Press et al., [2021]) introduces heuristics to design such attention bias, achieving early success in long-context extrapolation tasks for transformer architecture. Besides, Kerple (Chi et al., [2022]) and Sandwich building on the prior works, introduce trainable parameters to construct the attention bias matrix or construct attention bias through sinusoidal properties of position encoding.
On the other hand, another lively line of work delves into adjusting the RoPE. Inspired by the Neural Tangent Kernel (NTK) theory, NTK-aware Scaled RoPE (Xiong et al., [2023]) modifies the base parameter of the RoPE; LEX (Sun et al., [2022]) and PermuteFormer (Chen, [2021]) adds an exponential decay term; Positional Interpolation places linear scaling to every token; Dynamic-NTK (Huang and Yau, [2019]) gradually increases the scaling ratio. In the LLM era, YaRN (Peng et al., [2023]) linearly scales query and key with a temperature factor. Giraffe (Pal et al., [2023]) finds that high-frequency terms are hurt by the low-frequency terms that are usually ill-trained and proposes a scaling mechanism based on the power law to protect the well-trained high-frequency information.